OpenStudy (anonymous):

Solve the system using substitution 2x – y = 0 x + y = 6

OpenStudy (anonymous):

x=-6 . unless this is beginning algebra

OpenStudy (anonymous):

wait nevermind i read it wrong

hero (hero):

2x + x - y + y = 0 + 6 3x + 0 = 6 3x = 6 x = 6/2 x = 3

hero (hero):

y = 6